OpenStudy (anonymous):

Justin deposits $8,000 in a one-year CD at 3.1% interest, compounded daily. What is Justin's annual percentage yield (APY) to the nearest hundredth of a percent? (Hint: Find the interest using the compounding formula. Then use the simple interest formula to find the rate, as if the interest you found at first were simple interest.)

OpenStudy (anonymous):

What is the formula?

OpenStudy (wolf1728):

The formula for daily interest is Daily Interest = [(1+rate/365)^365]-1 And the rate would be entered as .031 Daily Interest = (1.031/365)^365 Daily Interest = 1.03148414607725-1 Daily Interest = .03148414607725 Daily Interest = 3.148414607725%